### v2

# GD doesn’t make the cut: Three ways that non-differentiability affects neural network training

Abstract
This paper investigates the distinctions between gradient methods applied to non-differentiable functions (NGDMs) and classical gradient descents (GDs) designed for differentiable functions. First, we demonstrate significant differences in the convergence properties of NGDMs compared to GDs, challenging the applicability of the extensive neural network convergence literature based on $L-smoothness$ to non-smooth neural networks. Next, we demonstrate the paradoxical nature of NGDM solutions for $L_{1}$-regularized problems, showing that increasing the regularization penalty leads to an increase in the $L_{1}$ norm of optimal solutions in NGDMs. Consequently, we show that widely adopted $L_{1}$ penalization-based techniques for network pruning do not yield expected results. Finally, we explore the Edge of Stability phenomenon, indicating its inapplicability even to Lipschitz continuous convex differentiable functions, leaving its relevance to non-convex non-differentiable neural networks inconclusive. Our analysis exposes misguided interpretations of NGDMs in widely referenced papers and texts due to an overreliance on strong smoothness assumptions, emphasizing the necessity for a nuanced understanding of foundational assumptions in the analysis of these systems.
Machine Learning, ICML
Gradient Descent (GD) and its variants (Duchi et al., 2011; Kingma & Ba, 2014; Lydia & Francis, 2019; McMahan & Streeter, 2010; Shi & Li, 2021; Tieleman et al., 2012; Zhang, 2018) have played a pivotal role in advancing image and language processing over the last two decades. These algorithms have found widespread use in neural network training because the are easy to implement, highly scalable, and iteratively approach stationary points, where the loss function’s gradient is zero (Bertsekas, 1997).
Despite being initially designed for differentiable loss functions, GDs are widely used to minimize the loss functions of non-differentiable neural networks. The loss functions of these non-differentiable neural networks takes a special form – they are differentiable almost everywhere. Consequently, when gradient descents are applied to these non-differentiable neural networks (which we refer to as non-differentiable gradient methods or NGDMs), they seldom encounter non-differentiable points during the training process. This trait, combined with recent findings that NGDMs converge to stationary points under mild regularity conditions (Bolte & Pauwels, 2021; Davis et al., 2020), has led standard references in the field to assume that non-differentiability in neural networks is inconsequential (see Section 6.1.5 of (Stevens et al., 2020), for instance).
Moreover, this assumption has sparked renewed interest in the dynamics of gradient descents amid complex loss landscapes (Ahn et al., 2022b; Dauphin et al., 2014; Reddi et al., 2019), with fewer studies explicitly addressing non-differentiability (Bertoin et al., 2021; Bolte & Pauwels, 2021; Davis et al., 2020). Additionally, it has led theoretical papers to develop results for continuously differentiable functions and then claim universality by running simulations on non-differentiable neural networks (Experiment 6 in Ahn et al. (2022b), Section 3.1 in Ma et al. (2022), Section 6 in Zhang et al. (2022) to state a few).
In our paper, we challenge this assumption, demonstrating that it leads to inaccuracies in several highly referenced papers and texts regarding training dynamics in neural networks. We analyze three different aspects of the training process and make the following contributions:
  • Convergence Analysis of ReLU Networks: We show that the rate of convergence of NGDMs is slower than that of GDs. Our findings strongly suggest that the many claims about rate of convergence made under the assumption of $L-smoothness$ have little bearing on the properties of NGDMs.
  • Solution to the LASSO problem: We show that NGDMs do not produce sparse solutions for the LASSO problem even for the simplest case of the $L_{1}$ penalized linear model. Furthermore, we show that the NGDM solution to the LASSO problem can produce a counterintutive result – increasing the LASSO penalty can result in a solution with a larger $L_{1}$ norm.
  • The Edge of Stability: We demonstrate that the edge of stability conjecture does not hold, even for gradient descents on all convex differentiable neural networks. Furthermore, we provide a counter-example to Ma et al. (2022)’s claim that subquadratic behavior around the local minimum is responsible for the Edge of Stability phenomenon.
In a broader context, our analysis emphasizes the critical role the assumptions we make play on our understanding of these complex systems. For NGDMs in particular, our results suggest a need to treat GDs and NGDMs as distinct systems with different dynamics, challenging the common assumption of their similarity.
1 Preliminaries
In the three sections that follow, we show how disregarding the non-differentiability of the loss function severely limits our understanding of different aspects of training dynamics. Each of these sections is largely self-contained, complete with a literature review, problem setup, and analysis. Throughout this paper, experiments involving deep networks consistently employ convolutional neural network (CNN) architectures akin to VGG16 (Simonyan & Zisserman, 2014). This choice is motivated by the architecture’s widespread usage in the papers referenced throughout our discussions. The present section addresses background and terminology that is pertinent to all the subsequent sections.
NDGM v/s GD:
The crux of our analyses lies in the structural differences between GDs and NDGMs. For a continuously differentiable loss function, $f(\beta)$, the well-studied GD sequence (Bertsekas, 1997; Boyd et al., 2004) is described by the recursion
$\beta_{t+1}=\beta_{t}-\alpha\nabla f(\beta_{t}),$ (1)
where $\alpha$ is the learning rate, and $\beta_{0}$ is a randomly initialized starting point.
For a continuous non-differentiable loss function, $g(\beta)$, the NDGM recursion is given by
$\tilde{\beta}_{t+1}=\tilde{\beta}_{t}-\alpha\widetilde{\nabla}g(\tilde{\beta}_{t}),$ (2)
where ~$\widetilde{\nabla}g(\tilde{\beta}_{t})=\nabla g(\tilde{\beta}_{t})$ if the gradient exists at $\tilde{\beta}_{t}$, else it is a heuristic measure  [^1] https://rb.gy/g74av. The deep learning literature widely assumes that the dynamics described by the recursions in (1) and (2) are nearly the same.
Convex non-differentiable loss functions:
In this paper, we show the problems with this assumption using a variety of convex loss functions. The majority of our analysis is performed on the penalized single layer neural network regression problem with a non-positive response vector, which Kumar (2023) has shown to be convex. The loss function for this problem is given by
$\begin{split}L_{1}(\beta;{\bf{Z}},{\bf{y}},\lambda_{1},\lambda_{2})&=||{\bf{y}}-\max(0,{\bf{Z}}\beta)||^{2}\\ &=+\lambda_{1}||\beta||_{1}+\lambda_{2}||\beta||_{2}^{2},\end{split}$ (3)
where $\beta$ is a $P\times 1$ parameter vector, ${\bf{Z}}$ is an $N\times P$ data matrix whose $i^{th}$ column is given by $z_{i}$, ${\bf{y}}$ is a $N\times 1$ non-positive response vector, $\lambda_{1}\geq 0$ is the LASSO penalty, and $\lambda_{2}\geq 0$ is the ridge penalty. Throughout this paper, we denote the learning rate by $\alpha$, and the $i^{th}$ entry in the parameter vector after $k$ iterations is $\beta_{k}[i]$, with $10\leq i\leq P-$ 1.
The subgradient method:
Our analysis heavily relies on the observation that for a convex non-differentiable loss functions, the NDGM recursion described in (2) effectively implements the subgradient method (Bertsekas, 1997; Shor, 2012). The theoretical properties of this method are well studied in the optimization literature, with a concise overview of key findings available in the lecture notes of Boyd et al. (2003) and Tibshirani (2015). The subgradient method and GD recursions have different properties, and these differences play a key role in our subsequent discussions. We reference the relevant portions from the lecture notes as they arise.
Constant, diminishing, and reducing step sizes:
Our formulations in equations (1) and (2) employ a constant learning rate. This choice bears some explaining, for which we first recall some definitions from the optimization literature (see Chapter 1 of (Bertsekas, 1997) for details). In the GD literature, the training regimen described in equation (1) is called a constant step-size regime. In this regime, the GD is initialized using some randomized approach, and iterated with a constant learning rate until the loss stops decreasing. Although constant step-size GDs approach the stationary point, they rarely converge to it except in special cases.
To ensure provable convergence to the stationary point, theoretical studies often adopt the diminishing step-size regime. This regime is characterized by the iteration
$\beta_{t+1}=\beta_{t}-\alpha_{t}\nabla f(\beta_{t}),$ (4)
where $\lim_{t\rightarrow\infty}\alpha_{t}=0$, and $\sum_{t}\alpha_{t}\rightarrow\infty$. Convergence results in this regime are derived under ideal conditions, and provide guarantees that the recursion will reach the stationary point at least in the theoretical limit.
In practice, diminishing step-size GDs are implemented with a stopping criteria which determines when the solution produced by the GD is good enough to end training. The combination of the diminishing step-size regime and stopping criteria commonly used in the neural network literature is what we term the “reducing” step-size regime. In this regime, training unfolds as follows: commencing with a weight of $\beta_{0}$, GD is run for some $N_{0}$ iterations with a learning rate of $\alpha_{0}$, followed by $N_{1}$ iterations with a reduced learning rate of $\alpha_{1}<\alpha_{0}$, and so on, until the engineer is satisfied with the solution. Essentially, the reducing step-size regime involves chaining together a finite number (say $T$) of constant step-size GD runs, with the $(i+1)^{th}$ run having an initialization of $\beta_{u}$ and a learning rate of $\alpha_{i}$, where $u=\sum_{j=0}^{i-1}N_{j}$. In the reducing step size regime, the theoretical guarantees described in the diminishing step-size regime may not be realized. Nevertheless, the limiting dynamics of the sequence in this regime are exactly those of a constant step-size GD, with learning rates and weight initializations chosen corresponding to the values at the beginning of the $T^{th}$ run. Accordingly, in this paper, we exclusively study the constant step-size problem, assuming that the weights and learning rates have been appropriately chosen.
2 NDGM and convergence analysis
Recent theoretical work by Davis et al. (2020) has shown that in the diminishing step-size regime, NDGMs converge to a local optimal solution in the theoretical limit under fairly general conditions. Unfortunately, these regularity conditions are rarely satisfied in practice. To illustrate this point, consider the Assumption A.3 in Davis et al. (2020), which requires that in the diminishing step size regime, the sum of the learning rates should diverge. As was previously stated, this is a standard assumption even in the thoeretical convergence analysis of GDs. However, the prevelant step-size reduction we observed in most training scripts involves decaying the learning rate by a constant multiplicative factor at predefined intervals (StepLR in PyTorch [^2] https://rb.gy/pub4s). With this decay scheme, the learning rates form a convergent geometric series, thus violating the assumption.
For the differentiable case, we know that even when the learning rates form a convergent geometric series, the GD sequence will come within some vicinity of the optimal solution. Therefore, the question arises if the same holds true for NDGM sequences. More precisely, can we state that the dynamics of NDGM and GD sequences are identical?
NDGM and GD dynamics are not identical:
The dynamics of the two sequences are not identical, and we demonstrate this by arriving at a contradiction. Suppose the sequence $\{\beta_{k}\}_{k=0}^{\infty}$ obtained by running NDGM on $L_{2}(\beta;{\bf{Z}}_{1},{\bf{q}}_{1})=L_{1}(\beta;{\bf{Z}}_{1},{\bf{q}}_{1},0,0.01)$ has the same properties as a GD sequence. Then, the entire NDGM sequence can be characterized using the ”Capture Theorem” as described below
Proposition 2.1.
Let $x_{k+1}=x_{k}-\alpha\nabla f(x_{k})$ be a sequence generated by running GD on a continuously differentiable function, $f(x)$ such that $f(x_{k+1})\leq f(x_{k})$. If $x^{*}$ is an isolated local minimum of $f(x)$ in a unit ball around $x^{*}$, and $||x_{0}-x^{*}||<1$, then $||x_{k}-x^{*}||<1$ for all $k>0$.
Proof.
Result follows by using $\bar{\epsilon}=1$ in Proposition 1.2.5 of (Bertsekas, 1997). ∎
Therefore, if $||\beta_{0}||_{\infty}<1$, and $L_{2}(\beta_{k+1};{\bf{Z}}_{1},{\bf{q}}_{1})\leq L_{2}(\beta_{k};{\bf{Z}}_{1},{\bf{q}}_{1})$ for all $k\geq 0$, then we expect $||\beta_{k}||_{\infty}<1$ for all $k\geq 0$. To validate this claim, we first generate a $20\times 500$ matrix ${\bf{Z}}_{1}$ with entries uniformly sampled on $[-1,1]$. Next, we generate a $20\times 1$ vector ${\bf{q}_{1}}$ with entries uniformly sampled on $[-1,0]$. Finally, we generate the sequence $\{\beta_{t}\}_{t=0}^{\infty}$, by running NDGM on $L_{2}(\beta;{\bf{Z}}_{1},{\bf{q}}_{1})$ with the entries in $\beta_{0}$ uniformly sampled on $[-1,1]$. As seen in Figure 1, although the loss function is a monotone decreasing function of the iteration number, the infinity norm of the NDGM sequence eventually takes values greater than 1, hence establishing the contradiction. The non-capturing described in Figure 1 is to be expected; as stated in Section 1 of Boyd et al. (2003), ”Unlike the ordinary gradient method, the subgradient method is not a descent method; the function value can (and often does) increase”.
Figure 1: Plot of the loss function ($L_{2}(\beta_{k};{\bf{Z}}_{1},{\bf{q}}_{1})$), and infinity norm ($||\beta_{k}||_{\infty}$) of the parameter vector as a function of the iteration number ($k$). Although the loss function is monotone decreasing, the infinity norm of the weights does not ”capture”, hence demonstrating that the dynamics of NDGM and GD are not identical
Convergence analysis and L-smoothness:
Given that the dynamics of GDs and NDGMs differ, the question arises what assumptions are appropriate for the study of convergence analysis in neural network training. Majority of the analyses in the literature (Défossez et al. (2020); Reddi et al. (2019); You et al. (2019); Zhang et al. (2022) to state a few) assume that the loss function, $f(\beta)$ is continuously differentiable, and is $L-smooth$, i.e., there exists a constant $L$, such that
$||\nabla f(x)-\nabla f(y)||\leq L||x-y||$ (5)
for all $x$, $y$ in the domain of $f$. However, this assumption breaks down for non-differentiable functions, where the gradient does not even exist at various points in the domain. This leads us to a pertinent inquiry: Can we salvage this assumption by confining its validity to a local context, possibly within a specific interval centered around the local optimal solution?
Unfortunately, even this cannot be done; as stated in (Laurent & Brecht, 2018), “local minima of ReLU networks are generically nondifferentiable. They cannot be waved away as a technicality, so any study of the loss surface of such network must invoke nonsmooth analysis”. Moreover, assuming that the loss function is $L-smooth$ grossly over-estimates the speed of convergence of NDGMs. To illustrate this, consider that in order to achieve an error level of $\epsilon$, GD applied to an $L-smooth$ convex loss function, $f$, would require $O(1/\epsilon)$ steps for convergence. In contrast, NDGM applied to a convex non-differentiable loss function, $g$, satisfying the Lipschitz condition
$||g(x)-g(y)||\leq L||x-y||$ (6)
would take $O(1/\epsilon^{2})$ steps (see Section 7.1.3 of Tibshirani (2015) for the result, and a visual illustration of how much slower $O(1/\epsilon^{2})$ is than $O(1/\epsilon)$). Thus, it is likely that the vast literature on the convergence analysis of GDs for loss functions satisfying (5) likely have little bearing on the dynamics of convergence in modern non-smooth neural networks. More generally, we expect the speed of convergence for smooth neural networks to be faster than that of non-smooth neural networks.
Experiments with NDGM convergence:
To validate our intuition in a more general context, we conduct experiments on CNNs of various depths (see the appendix for architecture details). For each network, we carry out two identical simulations, with the only difference being the activation function used. In the first simulation, we employ the non-smooth ReLU activation function, and in the second, we use the similarly shaped, but smooth and asymptotally equivalent, GELU activation function (Hendrycks & Gimpel, 2016).
As seen in Figure 2, the difference in convergence speeds between smooth and non-smooth networks widens as the depth of the neural network. This observation aligns with our expectations, given that increased depth introduces greater non-smoothness into the system. Hendrycks & Gimpel (2016) substantiate a similar trend of accelerated convergence in GELU networks compared to ReLU networks across diverse architectures and datasets (see Figures 2 through 7 in their paper).
Figure 2: Training curves comparing smooth and non-smooth CNNs (from left to right) of depth 11, 13, and 19 respectively. As the network depth increases, the non-smoothness in the system also increases. Consequently, the difference in convergence speeds between smooth and non-smooth networks widens with the increasing non-smoothness in the system.
Do results derived in the L-Smooth setting describe the behavior of all differentiable neural networks?:
In a subsequent section on the Edge of Stability, we show that insights based on the assumption of $L-smoothness$ are not universally applicable even across all Lipschitz continuous, convex, differentiable neural networks. We point out that $L-smoothness$ is one of the strongest assumptions which can be applied to a function after convexity. This assumption guarantees the existence of the first and second derivatives, and the boundedness of the second derivative—a collection of criteria that are unmet by the majority of the loss functions encountered in the neural network literature.
The way forward:
In this section, we show that NDGMs are slower to converge than GDs. Fortunately, most recent explorations into image and language models primarily leverage the transformer architecture (Vaswani et al., 2017) which employs the differentiable GELU activation function (Hendrycks & Gimpel, 2016). As a consequence, future iterations of neural networks may predominantly employ pure GDs. Therefore, it is crucial for the research community to establish a consensus on a set of assumptions that accurately represent these complex systems before meaningful progress can be achieved.
3 NDGM and the LASSO penalty
Zero weights play a critical role in model compression (Blalock et al., 2020; Han et al., 2015a, b; Li et al., 2016), and it is desirable to force weights down to zero whenever we can do so without sacrificing too much accuracy. A common approach to induce sparsity is to penalize the $L_{1}$ norm of the model weights, with the penalizing factor commonly referred to as the LASSO penalty. In the statistics literature, loss functions involving penalized $L_{1}$ norms are typically minimized using specialized algorithms designed for different use cases (Efron et al., 2004; Friedman et al., 2008; Mazumder & Hastie, 2012; Tibshirani, 1996; Tibshirani et al., 2005; Zou & Hastie, 2005). However when training neural networks, the penalized $L_{1}$ norm is used with arbitrary non-convex functions, with the understanding that running NDGM on the problem results in a near sparse optimal solution (Bengio, 2012; Goodfellow et al., 2016). As stated in Scardapane et al. (2017), ”Even if (the LASSO) has a non differentiable point in 0, in practice this rarely causes problems to standard first- order optimizers. In fact, it is common to simultaneously impose both weight-level sparsity with the $L_{1}$ norm, and weight minimization using the $L_{2}$ norm, resulting in the so-called elastic net penalization”. This expectation is not being met in reality, and there is strong evidence of users struggling with the interpretation of their LASSO results in the PyTorch forums  [^3] https://rb.gy/kh0qr  [^4] https://rb.gy/2vsry  [^5] https://rb.gy/6dfli. Given the wide variety of critical applications in which the LASSO finds use (Ghosh et al. (2021); Li et al. (2011); Li & Sillanpää (2012); Ogutu et al. (2012) to state a few), it is critical to understand the reliability of the NDGM solution for the LASSO problem.
NDGM solutions for the LASSO problem are not reliable:
We show that NDGM produces unreliable solutions even for the simple case of the LASSO penalized linear model. In particular, we show that increasing the LASSO penalty can result in a NDGM solution with a larger $L_{1}$ norm, completely defeating the purpose of the penalty. We first analytically demonstrate the unreliability for the simplest LASSO problem,
$L_{3}(\beta)=L_{1}(\beta;0,0,\lambda_{1},0)=\lambda_{1}||\beta||_{1},$ (7)
with $\lambda_{1}>0$, and then demonstrate the same issues for the general LASSO problem through simulations. We begin by showing that for any non-zero learning rate, $\alpha$, the NDGM sequence for (7) will be non-sparse with probability 1.
Proposition 3.1.
Let $\{\beta_{t}\}_{t=0}^{\infty}$ denote the sequence generated by running NDGM on (7), with a constant learning rate, $\alpha$, and the entries in $\beta_{0}$ being uniformly sampled from $[-1,1]$. Then $\beta_{k}$ will have all non-zero entries with probability 1 for all $k>0$.
Proof.
The NDGM iteration for (7) is described by the recursion
$\beta_{i}[k]=\begin{cases}\beta_{i-1}[k]-\alpha\lambda_{1}&\text{ if $\beta_{i-1}[k]>0$}\\ \beta_{i-1}[k]+\alpha\lambda_{1}&\text{ if $\beta_{i-1}[k]<0$}\\ 0&\text{ if $\beta_{i-1}[k]=0$}\end{cases}.$ (8)
From (8), it is clear that if the $k^{th}$ entry in the vector becomes zero during any iteration, then it stays zero for all subsequent iterations. Accordingly, suppose the $k^{th}$ entry in the NDGM sequence vector becomes 0 for the first time after $N$ iterations. Then we have
$\beta_{N}[k]=\beta_{0}[k]+(N-2m)\alpha\lambda_{1}=0,$ (9)
where $m$ is the number of times the $k^{th}$ entry of the parameter exceeds 0 in the first $N$ iterations. (9) can only hold if $\beta_{0}[k]$, is an integer multiple of $\alpha\lambda_{1}$. Since the feasible values are countable, and the set of initializations is uncountable, the probability of the occurrence has measure 0. ∎
In the next proposition, we show that the sequence does not converge, but oscillates between two fixed points.
Proposition 3.2.
The sequence $\{\beta_{t}\}_{t=0}^{\infty}$ described in the previous proposition does not converge. Furthermore, there is an integer $N_{0}$ such that for every $N\geq N_{0}$, the sequence oscillates between two vectors : $\gamma_{1}$ and $\gamma_{2}$, where $||\gamma_{1}||_{\infty}<\alpha\lambda_{1}$, and $||\gamma_{2}||_{\infty}<\alpha\lambda_{1}$.
Proof.
We prove the results assuming $\beta_{0}[k]>0$; the proof is similar when $\beta_{0}[k]<0$. From Proposition 3.1, we know that for any $N>0$, $\beta_{N}[k]\neq 0$ with probability 1. Therefore, (8) implies that starting from any $\beta_{0}[k]>0$, the sequence will decrease monotonically till it reaches a value between 00 and $\alpha\lambda_{1}$. Accordingly, let $\beta_{n_{k}}[k]=\gamma_{k}$ for some $n_{k}\geq 0$, with $\gamma_{k}\in(0,\alpha\lambda_{1})$. Since $\gamma_{k}-\alpha\lambda_{1}<0$, completing the recursion in (8) gives
$\beta_{n_{k}+m}[k]=\begin{cases}\gamma_{k}-\alpha\lambda_{1}&\text{ if m is odd}\\ \gamma_{k}&\text{ if m is even}.\end{cases}$ (10)
The result follows by choosing $N_{0}=\text{max}(n_{0},n_{1}\ldots n_{P-1})$. ∎
Equation (10) suggests the surprising result that with the same learning rate and initialization, a larger value of $\lambda_{1}$ can result in a NDGM solution with a larger $L_{1}$ norm, which completely defeats the purpose of the LASSO penalty. To better understand the issue, consider the 2D case, with the weights initialized as $\beta_{0}=[0.5053,0.5053]$. With $\lambda_{1}=1$ and $\alpha=0.01$, Proposition 3.2 states that the NDGM sequence will oscillate between $[0.0053,0.0053]$ and $[-0.0047,-0.0047]$, but with $\lambda_{1}=100$, and the same learning rate and starting point, the solution will oscillate between $[0.5053,0.5053]$ and $[-0.4947,-0.4947]$. Thus, at the point when NDGM appears to converge, the solution with $\lambda_{1}=1$ will have an $L_{1}$ norm of at most 0.0106, but solution with $\lambda_{1}=100$ case will have an $L_{1}$ norm of atleast 0.9894. Furthermore, note that reducing the step-size does not resolve this paradox. If, in both cases, we decrease the learning rate to $\alpha=0.001$ once the algorithm seems to have converged, then with $\lambda_{1}=1$, the NDGD sequence will oscillate between $[0.0003,0.0003]$ and $[-0.0007,-0.0007]$. However, with $\lambda_{1}=100$, the NDGM sequence will oscillate between $[0.0053,0.0053]$ and $[-0.0947,-0.0947]$, and the paradox persists.
To demonstrate the same issue in the broader setting, consider the loss function for the general LASSO problem given by
$L_{4}(\beta;{\bf{W}},{\bf{y}},\lambda_{1})=\frac{1}{N}||{\bf{y}}-{\bf{W}}\beta||_{2}^{2}+\lambda_{1}||\beta||_{1},$ (11)
where ${\bf{W}}$ is an arbitrary $20\times 500$ dimensional data matrix, and ${\bf{y}}$ is an arbitrary $20\times 1$ response, each of whose entries are sampled from the uniform distribution on $[-1,1]$. We run NDGM on (11) twice with the same learning rate and initialization, once with $\lambda_{1}=0.01$, and once with $\lambda_{1}=10$. At the end of the two runs, the optimal solution with $\lambda_{1}=0.1$ has an $L_{1}$ norm of 1.62, and the optimal solution with $\lambda_{1}=10$ has an $L_{1}$ norm of 25.4 i.e., a 1000 fold increase in the value of $\lambda_{1}$ results in a more than 15 fold increase in the $L_{1}$ norm of the optimal solution!
Connections to the subgradient method:
The counter-intuitive behavior seen in the example above can be explained using the subgradient method. Since the loss function described in equation (7) satisfies the Lipschitz condition (see equation (6)) with $L=\lambda_{1}$, we have from Boyd et al. (2003) (Section 2 on the Constant Step Size) that
$\lim_{k\rightarrow\infty}||\beta_{k}||_{1}<\alpha\lambda_{1}.$ (12)
Therefore, for the same value of $\alpha$, the larger value of $\lambda_{1}$ will result in a larger error in estimation, which is precisely what we are seeing above.
Diminishing v/s reducing step sizes:
This illustration also highlights the contrast between the diminishing and reducing step size regimes outlined in the preliminaries. Under the diminishing step-size regime, $\alpha\rightarrow 0$ in the theoretical limit, and therefore, Proposition 3.1 guarantees that the optimal solution for (7) will be 0 for all $\lambda_{1}>0$. On the other hand, in the reducing step-size regime, $\alpha>0$ at the end of training, and therefore, the error in estimation becomes a critical determinant of the optimal value.
Experiments with deep networks:
In their seminal work on network pruning, Han et al. (2015b) propose the utilization of $L_{1}$ regularization to encourage weights to approach 0, facilitating the pruning process. This approach is widely adopted in the literature (Wen et al., 2016; Xiao et al., 2019; Yang et al., 2019), and is explicitly endorsed in NVIDIA’s guides for training standard object detection models such as RetinaNet and YOLOv4666See the recommendation of the regularizer in the Training config section of https://rb.gy/20y13m and https://rb.gy/i1sb6s. To challenge the general applicability of this suggestion, we fine-tune two pretrained VGG16 models using the $L_{1}$-penalized cross entropy loss with a learning rate of $\alpha=0.01$, with one model using a regularization parameter of $\lambda_{1}=0.001$ and the other with $\lambda_{1}=0.1$. At the end of fine-tuning, the $L_{1}$ norm of the weights in the model fine-tuned with $\lambda_{1}=0.1$ is more than 30 times larger (69178.44) than that of the model fine-tuned with $\lambda_{1}=0.001$ (2113.83). Furthermore, the latter model has 99.7% of its weights less than $10^{-5}$, whereas the former model has only 0.01% of its weights less than $10^{-5}$, further underscoring the discrepency. We conducted experiments with different values of $\alpha$, and the widely used Adam optimizer (Kingma & Ba, 2014), and in each case, we observed the same paradox. Our code is provided in the appendices.
NDGM and the way forward with sparsity:
While NDGMs do not inherently generate sparse solutions, iterative methods have been developed to produce sparse solutions for the $L_{1}$ regularized linear problem (see Table 1 in Bottou (2010)), and $L_{1}$ regularized neural networks in general (Ma et al., 2019; Siegel et al., 2020). The realization that NDGMs lack the capability to produce sparse solutions predates the widespread adoption of deep learning. As mentioned in (Xiao, 2009), “… stochastic gradient descent (SGD), has limited capability of exploiting problem structure in solving regularized learning problems”. Despite this, a prevalent misconception persists in the deep learning literature, erroneously suggesting that such algorithms yield sparse solutions, and that the solutions generated by deep learning frameworks like PyTorch (Paszke et al., 2017) ought to be comparable to those generated using traditional machines learning frameworks like glmnet (Friedman et al., 2021) or scikit-learn (Kramer & Kramer, 2016). We point out that there exits a fundamental disparity between traditional machine learning frameworks like glmnet and deep learning frameworks like PyTorch. glmnet focuses on a narrow set of optimization problems, concealing implementation intricacies while guaranteeing optimality. Conversely, PyTorch provides greater flexibility in constructing models but lacks similar guarantees, placing the responsibility for solution quality on users. Attaining a solution at a non-differentiable point poses a challenging problem, underscoring the significance of exploring related studies and tools rather than opting for a simplistic implementation, particularly when precision is paramount.
Connections to the Edge of Stability:
Proposition 3.2 shows that the NDGM sequence for the non-differentiable, Lipschitz continuous, convex loss function described in (7) will not diverge to $\infty$ for any finite value of $\alpha$. This is a special case of unstable convergence, a topic we cover in greater detail in the next section.
4 NDGM and the Edge of Stability
In recent years, there has been a growing interest on the topic of “unstable convergence”. The literature on the subject (Ahn et al., 2022a, b; Arora et al., 2022; Chen & Bruna, 2022; Cohen et al., 2021; Li et al., 2022) is motivated by the finding that unlike convex quadratic forms, gradient descent on neural networks do not diverge to $\infty$ even when the learning rate is greater than $\alpha^{*}=2/\eta$, where $\eta$ is the dominant eigenvalue of the loss function’s Hessian. Instead, the loss function has been empirically demonstrated to converge “unstably”, reducing non-montonically on the long run. $\alpha=\alpha^{*}$ is referred to in the literature as the ”Edge of Stability”, since it separates regions of ”stable convergence” ($\alpha<\alpha^{*}$) from regions of “unstable convergence” [^7] We define unstable convergence to mean “non-divergence to $\infty$”. Our definition is adapted from Definition 1.1 of Arora et al. (2022). However, we deviate from their definition due to its dependence on the existence of the Hessian at all points between the current and next iterate, a condition that need not be satisfied in non-smooth neural networks. ($\alpha>\alpha^{*}$).
The derivation of the Edge of Stability condition assumes that the loss function, $f$, is $L-smooth$ (see equation (5)). Since the result has been demonstrated in most convex optimization settings (Jacot et al., 2018; Li & Liang, 2018), it is conjectured that the result ought to hold even in the case of non-convex, non-differentiable neural networks. This conjecture is incorrect even for Lipschitz continuous, differentiable, convex functions satisfying (6) as shown in the proposition below.
Proposition 4.1.
All convex non-smooth loss functions having bounded subgradients, or satisfying equation (6) will show unstable convergence.
Proof.
Note that any loss function satisfying (6) has subgradients bounded above by $L$. Additionally, as outlined in Section 1 of Boyd et al. (2003), the deductions derived from the subgradient method hold equivalently for differentiable functions. This equivalence arises from the fact that the only admissible value for the subgradient of a differentiable convex function at any given point is, indeed, its gradient. Accordingly, using the convergence properties of the subgradient methods with constant step sizes from Section 2 of Boyd et al. (2003), we have that $\lim_{k\rightarrow\infty}f(x_{k})-f^{*}\leq\alpha L^{2}<\infty$, hence the result. ∎
Figure 3: Training curves for $\alpha=0.1$, $\alpha=1$, and $\alpha=10$ with the Lipschitz continuous Huber loss function. As the learning rate increases, the loss starts oscillating with larger frequencies, but never diverges to $\infty$. This is to be expected from Proposition 4.1, since the Huber loss function is convex, non-smooth and has bounded subgradients.
In the previous section, we analytically derive the unstable convergence for the LASSO problem described in (7). An example of a differentiable non-smooth convex function satisfying Proposition 4.1 is the Huber loss function, which finds use in many state of the art object detection models (Girshick, 2015; Liu et al., 2016; Ren et al., 2015). The Huber loss function is non-smooth because it is once but not twice differentiable. For a regression problem involving an arbitrary $50\times 1$ response vector, $\mathbf{y}=[y_{1},y_{2}\ldots y_{N}]^{T}$, and an arbitrary $50\times 200$ data matrix, $\mathbf{Z}$, the Huber loss function is defined as
$L_{7}(\beta)$ $=\frac{1}{N}\sum l(i),\quad\text{where}$ (13)
$l(i)$ =$=\begin{cases}\frac{1}{2}(y_{i}-z_{i}^{T}\beta)^{2}&\text{if }\left|(y_{i}-z_{i}^{T}\beta)\right|<1\\ \left(|y_{i}-z_{i}^{T}\beta|-\frac{1}{2}\right)&\text{otherwise}\end{cases}$
where $z_{i}$ is the $i^{th}$ column of $\mathbf{Z}$. Figure 3 shows that the NDGM sequence for (13) does not diverge toward $\infty$, even with a high learning rate, such as $\alpha=10$, which aligns with our expectations. Furthermore, the figure reveals an intriguing loss pattern when $\alpha=1$: the loss initially increases before the oscillations become noticeable. This peculiar behavior has been documented in the literature (Ahn et al., 2022b; Arora et al., 2022), although a conclusive explanation for this phenomenon remains elusive at this time.
Proposition 4.1 suggest that it is unclear and the examples that follow show that the common approach of extrapolating results developed for $L-smooth$ loss functions to the non-smooth or non-differentiable case is problematic. In the next subsection, we demonstrate the problems with this approach by providing an example of an extrapolation which breaks down in the non-smooth case.
Subquadratic losses and unstable convergence
In section 3.1 of Ma et al. (2022), the authors employ second-order finite differences to investigate the curvature of VGG16 and ResNet in the vicinity of the local minimum. They conclude that subquadratic growth of the loss function around this minimum point explains the edge of stability phenomenon. However, their analysis is flawed because VGG16 and ResNet, as ReLU networks, have local minima that are generically non-differentiable (Laurent & Brecht, 2018). As a result, the points of interest lie where their approximation breaks down.
To challenge the broad applicability of their statement, consider the Huber loss function depicted in Equation (13). This function is overparameterized and therefore possesses a global minimum of 0. Consequently, within the vicinity of this global minimum, the loss function exhibits quadratic growth. In such instances, Ma et al. (2022)’s result states that we will not see unstable convergence. However, contrary to their suggestions, Proposition 4 guarantees unstable convergence under these circumstances, as evidenced in Figure 3.
The discrepancy arises because the Huber loss function lacks the twice differentiability required by Definition 4.1 of Ma et al. (2022)’s paper. In contrast, all $L-smooth$ functions will exhibit subquadratic growth near their local minimum during unstable convergence, as noted by (Ma et al., 2022). This occurs because $L-smooth$ functions are twice differentiable by definition. This example highlights the challenges of extending results developed under strong assumptions like $L-smoothness$ to the general case.
The way forward:
We demonstrate that our understanding of the Edge of Stability will be incomplete until we more precisely quantify conditions under which we expect it to occur. Furthermore, considering ReLU networks’ non-differentiable minima, without further research into NDGM behavior around non-differentiable points our grasp of training dynamics lacks completeness. Consequently, we propose employing the convex non-differentiable single-layer neural network presented in (Kumar, 2023) as the toy problem for theory development. Studying this neural network offers several advantages: 1) its convex nature renders it more amenable to analysis compared to non-convex functions, 2) we possess analytical knowledge of the global optimum, and 3) the global optimum exists at a point of non-differentiability—the setup which is particularly relevant to ReLU networks.
5 Conclusion
In this paper, we highlight how non-differentiability alters training dynamics across three domains: convergence analysis, the LASSO problem, and the Edge of Stability. In each case, we show that problems arise because of incorrect assumptions made about the contour of the loss surface. Our results suggest the critical need to reach consensus on a set of assumptions which can be made about the loss surface before progress can be made.
Appendix A Appendix A - Architecture and code for Figure 2
The code to construct our plots in Figure 2, was taken from https://github.com/kuangliu/pytorch-cifar/blob/master/models/vgg.py, with three minor modifications. First, we removed the batch normalization as part of each layer. We did this because, while it is known that batch normalization alters the training dynamics (Ioffe & Szegedy, 2015), and our work is interested in the effect of network smoothness on convergence speed in the absence of confounders. Second, we replaced max pool with average pool to ensure differentiability of the layer, and third, we used a simpler layer configuration than one provided in the link. The configuration we used is provided below:
cfg = {
    ’VGG11’: [16, ’M’, 32, ’M’, 32, 32, ’M’, 64, 64, ’M’, 64, 64, ’M’],
    ’VGG13’: [16, 16, ’M’, 32, 32, ’M’, 32, 32, ’M’, 64, 64, ’M’, 64, 64, ’M’],
    ’VGG19’: [16, 16, ’M’, 32, 32, ’M’,
              32, 32, 32, 32, ’M’,
              64, 64, 64, 64, ’M’,
              64, 64, 64, 64, ’M’],
Appendix B Appendix B - Code for paradoxes relating to the LASSO
The training code was generated using ChatGPT and verified for correctness. The code below utilizes SGD with a learning rate of 0.01, though can trivially be changed to any other.
import torch
import torch.nn as nn
import torch.optim as optim
from torchvision import datasets, transforms, models
# Set random seed for reproducibility
seed = 42
torch.manual_seed(seed)
torch.backends.cudnn.deterministic = True
torch.backends.cudnn.benchmark = False
import numpy as np
np.random.seed(seed)
# Set device to GPU if available
device = torch.device("cuda" if torch.cuda.is_available() else "cpu")
# Define transform and download CIFAR-10 dataset
transform = transforms.Compose([
    transforms.ToTensor(),
    transforms.Normalize((0.5, 0.5, 0.5), (0.5, 0.5, 0.5)),
train_dataset = datasets.CIFAR10(root=’./data’, train=True, download=True, transform=transform)
train_loader = torch.utils.data.DataLoader(train_dataset, batch_size=512, shuffle=True, num_workers=4)
# Load pretrained VGG16 model
model = models.vgg16(pretrained=True)
model.to(device)
# Define L1 regularization strength
l1_strength = 0.1
# Define the loss function (CrossEntropyLoss + L1 penalty)
criterion = nn.CrossEntropyLoss()
optimizer = optim.SGD(model.parameters(), lr=0.01)
# Training loop
epochs = 150
for epoch in range(epochs):  # Adjust the number of epochs as needed
    for inputs, labels in train_loader:
        inputs, labels = inputs.to(device), labels.to(device)
        # Forward pass
        outputs = model(inputs)
        loss_ce = criterion(outputs, labels)
        # L1 penalty
        l1_penalty = 0
        for param in model.parameters():
            l1_penalty += torch.abs(param).sum()
        # Add L1 penalty to the loss
        loss = loss_ce + l1_strength * l1_penalty
        # Backward and optimize
        optimizer.zero_grad()
        loss.backward()
        optimizer.step()
    print(f’Epoch [{epoch+1}/100], L1 Penalty: {l1_penalty.item():.4f}’)
print("Training complete!")
total_params = 0
num_small_params = 0
threshold = 1e-5
for name, param in model.named_parameters():
    total_params += param.numel()
    if ’weight’ in name or ’bias’ in name:
        num_small_params += torch.sum(torch.abs(param) < threshold).item()
fraction_small_params = num_small_params / total_params
print(f’Fraction of parameters with values less than 10^-5: {fraction_small_params:.5f}’)